OpenStudy (anonymous):

Write a quadratic equation having the given numbers as solutions. -8 and -4

OpenStudy (anonymous):

(x+8) * (x+4) = x^2 + 12x + 32

OpenStudy (anonymous):

(x+8)(x+4)=0 x^2+12x+32=0

OpenStudy (anonymous):

Thank you both